Supposed the disposable income ranges from 1,110,000 to 1,290,000 and it increases by 30,000 per year. Mpc = .80




When disposable income is 1,000,000 what is the level of consumption is equal to?

1
Expert's answer
2022-02-21T15:02:02-0500

"\text{C}=\text{a}+\text{b}Y_d\\
\text{where;}\\
\text{a - autonomous consumption}\\
\text{b - MPC}\\
Y_d\, \text{- disposable income}\\
\text{therefore,}\\
C=30,000+(0.80*1,000,000)\\
C=30,000+800,000\\
C=830,000."